OpenStudy (rockycatsrule):

Ari’s chickens eat their pellets at a constant rate throughout each day. What is the slope of the line that represents the ounces of pellets eaten? Ari’s Chicken Pellets Days 2 4 6 8 Ounces of Pellets remaining 39 33 27 21 fans and metals plz help me

OpenStudy (hungergames13):

-3

OpenStudy (rockycatsrule):

thx so much

OpenStudy (rockycatsrule):

can you help on some others too

OpenStudy (hungergames13):

yup

OpenStudy (rockycatsrule):

ok

OpenStudy (alexandervonhumboldt2):

@hungergames13 please explain your answer. Giving direct answers is against the rules.

OpenStudy (rockycatsrule):

Carmen hikes at a constant rate. The graph shows the distance Carmen has traveled at different times during his hike. What is the slope of the line that represents the situation? Enter the slope as a fraction.

OpenStudy (rockycatsrule):

i will post the file

OpenStudy (rockycatsrule):

First quadrant grid, x-axis numbered 0 to 10 and labeled Hours, y-axis numbered 0 to 10 and labeled Miles. Line passes through labeled points at (3, 2) and (6, 4).

OpenStudy (hungergames13):

is there a graph?

OpenStudy (rockycatsrule):

yes just give me a min

OpenStudy (hungergames13):

ok

OpenStudy (rockycatsrule):

there it is

OpenStudy (hungergames13):

to find slope use rise/run count how many up then how many across from any given two points

OpenStudy (hungergames13):

m= 2/3

OpenStudy (rockycatsrule):

thx

OpenStudy (hungergames13):

YW

OpenStudy (rockycatsrule):

Juan always saves the same amount from his weekly allowance. The table shows how much he has saved at different times. Which equation represents the situation? Juan’s Savings Weeks 3 6 9 12 Amount saved 19 28 37 46 A. y – 3 = −3(x – 19) B. y – 28 = 3(x – 6) C. y + 9 = −3(x + 37) D. y + 46 = 3(x – 12)